Q1. Find mean deviation from the mean of the following data. Size of the items ?? 4 6 8 10 12 14 16 Frequency ?? 2 4 5 3 2 1 3 where mean is given 10 marks. Q2. The daily sale of chocolate in Othammarket given below: Monday Tuesday Wednesday Thursday Friday Saturday Sunday 140kg 50kg 60kg 70kg 80kg 90kg 100kg The average daily sale is 75kg. Calculate the variance and the standard deviation of the above data. Q3. The rank of some 8 students in two subjects X and Y are given below. ???? ? 5 2 8 1 6 7 3 4 ???? ? 4 5 1 3 8 6 2 7 Calculate the rank correlation coefficient. Q4.Calculate the coefficient correlation from the following data. ? 64 65 66 67 73 ? 45 50 55 60 70 Q5. If the probabilities that an automobile mechanic will servie 3, 4, 5, 6, 7 or 8 or more cars on any given workday are, respectively, 0.12, 0.19, 0.28, 0.24, 0.10, and 0.07, what is the probability that he will service at least 5 cars on his next day at work? Q6. What is the probability of getting a total of 7 or 11 when a pair of fair dice are tossed. Question Four (7+8=15 marks) Q7. If is believed that 89% of the population is right-handed. 150 People are randomly surveyed, are 135 of them claim to be righthanded. Calculate the ?-test statistic. Q8. Find the ?-test value and degree of freedom for the following two sets of data. ?? 2 5 7 9 11 13 ?? 1 2 3 4 5 6
Answer #1

Q1:

x f |x-10|*f
4 2 12
6 4 16
8 5 10
10 3 0
12 2 4
14 1 4
16 3 18
Total 20 64

Q5:  Probability that he will service at least 5 cars on his next day at work= 0.28+0.24+0.10+0.07=0.69

Q6: Define, A=event of getting a total of 7 or 11 when a pair of fair dice are tossed then

A={(1,6),(2,5),(3,4),(4,3),(5,2),(6,1),(5,6),(6,5)} and total no. of outcomes=62=36.

Probability of getting a total of 7 or 11=P(A)=n(A)/36=8/36=2/9

where, n(A)=no. of elements in A =8